# CaRtGS: Computational Alignment for Real-Time Gaussian Splatting SLAM

### v2

@@ III-B2 Adaptive Optimization
where $r_{k+1}^{0}$ is the initial training iteration count assigned to the new keyframe, and $l_{k+1}$ is its initial training loss value. We then select a keyframe $v^{\prime}$ randomly from the subset of keyframes with remaining iterations, defined as $\left\{v_{i}|r_{i}>0,\forall r_{i}\in\mathcal{R}_{k}\right\}$, to train the 3D Gaussians Map $\mathcal{G}$. Post-training, we decrement the training iteration count for the selected keyframe by one, adjusting $r^{\prime}$ to $r^{\prime}-1$, and also update the corresponding training loss value $l^{\prime}$.
@@
where $\overset{d_{k}}{\prod}(\cdot)$ donates top $d_{k}$ largest elements, $d_{k}=\max(1,\frac{k}{d})$, and $d$ is a hyperparameter. This method prioritizes keyframes with higher training loss values for the photorealistic rendering module, effectively tackling the long-tail optimization as demonstrated in LABEL:Fig:adaptive-opt.
III-B3 Opacity Regularization
In the typical application of 3DGS, the rendered loss $\mathcal{L}_{rendered}$ is utilized to refine the 3D Gaussian primitives [6]. To efficiently manage memory usage and model size, we have devised a strategy that encourages the elimination of Gaussians in areas where they do not contribute to the rendering process. Since the presence of a Gaussian is primarily indicated by its opacity $o$, we impose a regularization term $\mathcal{L}_{o}$ on this attribute. The complete formulation of our training loss $\mathcal{L}$ is as follows:
@@ IV Experiments
In this section, we present a comparative analysis of CaRtGS against state-of-the-art GS-SLAM systems. This evaluation spans multiple scenarios, including those captured using monocular and RGB-D cameras. Furthermore, we perform an ablation study to substantiate the efficacy of the novel techniques introduced in our approach.
@@ IV-A Setup
Dataset. We conducted evaluations on two distinct camera systems: monocular and RGB-D. These assessments were carried out on two renowned datasets: Replica [26] and TUM-RGBD [27]. To ensure data consistency, we employed a soft time synchronization to align the sensor data and ground truth with a precision of $\Delta t=0.08s$.
Implementation Detail. All experimental evaluations were conducted on a desktop with an Nvidia RTX 4090 GPU, an AMD Ryzen 9 7950X CPU, and 128 GB RAM. We implement our method on top of the Photo-SLAM [15]. Given the iterative nature of SLAM, we do not confine the optimization of photorealistic rendering to a predetermined number of iterations. Instead, the process continues dynamically until a termination signal is received from the front-end tracker, ensuring that the optimization is responsive to real-time tracking data. We retained most of the original hyperparameters from the 3DGS [6]. However, we increased the adaptive densification interval (500). Additionally, we raised the minimum gradient threshold for adaptive densification (0.001) and the minimum opacity threshold for pruning (0.02). By default, we set $d=4$ and $\lambda_{o}=0.001$. On Replica, we use $r_{k+1}^{0}=8$, whileas $r_{k+1}^{0}=2$ on TUM-RGBD.
Evaluation. We performed all experiments $5$ times to ensure statistical robustness and rendered original-resolution images for each estimated camera pose. To measure performance, we utilized the evo toolkit [^1] https://github.com/MichaelGrupp/evo and the torchmetrics toolkit [^2] https://github.com/Lightning-AI/torchmetrics. We recorded various performance indicators, including Absolute Trajectory Error (ATE) to assess the accuracy of localization, and metrics for evaluating photorealistic rendering such as Peak Signal-to-Noise Ratio (PSNR), Structural Similarity Index (SSIM), and Learned Perceptual Image Patch Similarity (LPIPS). Additionally, we reported the model size by showing the 3D Gaussian points. All performance indicators are reported in the format of mean $\pm$ standard deviation.

### v4

@@ III-C2 Adaptive Optimization
Table I: Quantitative Results on Replica.
Monocular Photo-SLAM [17] ATE $0.20\pm 0.02$ $2.95\pm 6.23$ $0.91\pm 0.39$ $0.11\pm 0.01$ $0.17\pm 0.00$ $0.15\pm 0.00$ $0.24\pm 0.04$ $0.10\pm 0.02$
FPS $36.91\pm 0.75$ $36.41\pm 0.66$ $34.48\pm 0.52$ $34.60\pm 0.36$ $35.98\pm 0.49$ $34.40\pm 0.29$ $36.37\pm 0.66$ $33.32\pm 0.28$
IPF $2.66\pm 0.11$ $2.31\pm 0.05$ $2.30\pm 0.06$ $2.29\pm 0.05$ $2.30\pm 0.04$ $2.03\pm 0.02$ $2.22\pm 0.02$ $2.30\pm 0.08$
PSNR $35.02\pm 0.45$ $32.75\pm 5.37$ $31.19\pm 0.65$ $31.13\pm 0.53$ $32.94\pm 0.18$ $28.74\pm 0.39$ $30.56\pm 0.38$ $31.69\pm 0.25$
Points $78.40\text{k}\pm 2.94\text{k}$ $97.04\text{k}\pm 31.16\text{k}$ $99.40\text{k}\pm 1.67\text{k}$ $76.36\text{k}\pm 3.19\text{k}$ $75.98\text{k}\pm 3.39\text{k}$ $0.11\text{m}\pm 6.27\text{k}$ $0.12\text{m}\pm 5.56\text{k}$ $81.10\text{k}\pm 1.81\text{k}$
Ours (Photo-SLAM) ATE $0.22\pm 0.06$ $2.97\pm 6.24$ $1.53\pm 1.37$ $0.12\pm 0.01$ $0.17\pm 0.01$ $0.17\pm 0.00$ $0.52\pm 0.48$ $0.09\pm 0.00$
FPS $36.65\pm 0.46$ $36.08\pm 0.47$ $33.90\pm 0.28$ $34.88\pm 0.68$ $35.96\pm 0.54$ $33.58\pm 0.20$ $36.65\pm 0.29$ $33.73\pm 0.26$
IPF $8.10\pm 0.21$ $7.76\pm 0.23$ $8.05\pm 0.13$ $7.15\pm 0.09$ $7.35\pm 0.13$ $7.40\pm 0.11$ $7.33\pm 0.24$ $7.67\pm 0.04$
PSNR $34.58\pm 0.31$ $34.97\pm 4.96$ $33.52\pm 0.12$ $33.26\pm 0.08$ $35.22\pm 0.23$ $31.92\pm 0.26$ $31.99\pm 1.15$ $34.39\pm 0.16$
Points $\boldsymbol{38.32\text{k}\pm 1.97\text{k}}$ $\boldsymbol{48.37\text{k}\pm 11.77\text{k}}$ $\boldsymbol{64.07\text{k}\pm 1.03\text{k}}$ $\boldsymbol{54.93\text{k}\pm 0.91\text{k}}$ $\boldsymbol{53.67\text{k}\pm 1.13\text{k}}$ $\boldsymbol{87.49\text{k}\pm 2.99\text{k}}$ $\boldsymbol{73.44\text{k}\pm 2.84\text{k}}$ $\boldsymbol{58.92\text{k}\pm 1.30\text{k}}$
RGBD Photo-SLAM [17] ATE $0.45\pm 0.05$ $0.35\pm 0.04$ $1.13\pm 0.14$ $0.37\pm 0.02$ $0.44\pm 0.05$ $0.30\pm 0.02$ $0.33\pm 0.04$ $0.18\pm 0.00$
FPS $31.61\pm 0.53$ $31.96\pm 0.32$ $30.43\pm 0.81$ $29.33\pm 0.52$ $27.87\pm 0.54$ $27.49\pm 0.52$ $29.87\pm 0.91$ $27.37\pm 0.52$
IPF $3.43\pm 0.09$ $3.04\pm 0.12$ $3.18\pm 0.04$ $3.28\pm 0.04$ $3.10\pm 0.05$ $3.17\pm 0.05$ $3.12\pm 0.05$ $3.20\pm 0.05$
PSNR $36.83\pm 0.32$ $36.79\pm 0.29$ $32.45\pm 0.38$ $33.38\pm 0.07$ $35.13\pm 0.39$ $30.13\pm 2.14$ $33.80\pm 0.36$ $34.53\pm 0.87$
Points $81.34\text{k}\pm 2.95\text{k}$ $79.24\text{k}\pm 1.71\text{k}$ $0.12\text{m}\pm 4.04\text{k}$ $93.03\text{k}\pm 3.79\text{k}$ $0.12\text{m}\pm 1.61\text{k}$ $0.19\text{m}\pm 2.70\text{k}$ $0.16\text{m}\pm 8.84\text{k}$ $0.14\text{m}\pm 2.09\text{k}$
Ours (Photo-SLAM) ATE $0.48\pm 0.04$ $0.38\pm 0.06$ $1.10\pm 0.19$ $0.38\pm 0.02$ $0.56\pm 0.10$ $0.31\pm 0.01$ $0.34\pm 0.03$ $0.18\pm 0.00$
FPS $30.84\pm 0.37$ $31.49\pm 0.31$ $30.04\pm 0.43$ $28.76\pm 0.58$ $28.64\pm 0.66$ $27.81\pm 0.62$ $29.55\pm 0.55$ $26.87\pm 0.31$
IPF $10.45\pm 0.30$ $9.90\pm 0.26$ $10.06\pm 0.21$ $10.40\pm 0.40$ $10.71\pm 0.35$ $9.95\pm 0.66$ $9.25\pm 0.40$ $9.97\pm 0.06$
PSNR $35.54\pm 0.28$ $37.74\pm 0.41$ $33.40\pm 0.29$ $33.84\pm 0.27$ $35.64\pm 0.41$ $29.38\pm 3.70$ $34.30\pm 0.64$ $36.54\pm 0.19$
Points $39.74\text{k}\pm 1.11\text{k}$ $54.61\text{k}\pm 2.58\text{k}$ $79.29\text{k}\pm 3.24\text{k}$ $68.03\text{k}\pm 2.06\text{k}$ $75.58\text{k}\pm 4.31\text{k}$ $0.11\text{m}\pm 3.74\text{k}$ $0.10\text{m}\pm 1.21\text{k}$ $0.10\text{m}\pm 2.72\text{k}$
GS-ICP-SLAM[16] ATE $0.19\pm 0.00$ $0.13\pm 0.00$ $0.18\pm 0.00$ $0.19\pm 0.01$ $0.22\pm 0.01$ $0.16\pm 0.00$ $0.16\pm 0.00$ $0.11\pm 0.01$
FPS $30.00\pm 0.00$ $30.00\pm 0.00$ $30.00\pm 0.00$ $30.00\pm 0.00$ $30.00\pm 0.00$ $30.00\pm 0.00$ $30.00\pm 0.00$ $30.00\pm 0.00$
IPF $2.88\pm 0.00$ $2.37\pm 0.01$ $2.88\pm 0.01$ $2.87\pm 0.00$ $2.91\pm 0.01$ $2.90\pm 0.07$ $2.84\pm 0.07$ $2.67\pm 0.01$
PSNR $40.57\pm 0.03$ $40.96\pm 0.11$ $32.77\pm 0.16$ $31.60\pm 0.07$ $38.84\pm 0.04$ $35.54\pm 0.06$ $37.81\pm 0.06$ $38.54\pm 0.05$
Points $1.57\text{m}\pm 0.85\text{k}$ $1.57\text{m}\pm 7.30\text{k}$ $1.54\text{m}\pm 2.51\text{k}$ $1.55\text{m}\pm 9.54\text{k}$ $1.60\text{m}\pm 10.33\text{k}$ $1.55\text{m}\pm 2.86\text{k}$ $1.55\text{m}\pm 0.70\text{k}$ $1.54\text{m}\pm 3.78\text{k}$
Ours (GS-ICP-SLAM) ATE $0.25\pm 0.14$ $0.12\pm 0.00$ $0.28\pm 0.10$ $0.19\pm 0.02$ $0.24\pm 0.01$ $0.16\pm 0.00$ $0.16\pm 0.00$ $0.11\pm 0.00$
FPS $30.00\pm 0.00$ $30.00\pm 0.00$ $30.00\pm 0.00$ $30.00\pm 0.00$ $30.00\pm 0.00$ $30.00\pm 0.00$ $30.00\pm 0.00$ $30.00\pm 0.00$
IPF $12.10\pm 0.07$ $12.36\pm 0.05$ $10.42\pm 0.08$ $10.97\pm 0.04$ $11.47\pm 0.05$ $11.03\pm 0.03$ $11.68\pm 0.08$ $10.46\pm 0.05$
PSNR $\boldsymbol{42.60\pm 0.05}$ $\boldsymbol{42.33\pm 0.03}$ $\boldsymbol{36.95\pm 0.42}$ $\boldsymbol{36.87\pm 0.03}$ $\boldsymbol{39.77\pm 0.08}$ $\boldsymbol{36.46\pm 0.10}$ $\boldsymbol{39.19\pm 0.05}$ $\boldsymbol{39.38\pm 0.23}$
Points $0.76\text{m}\pm 7.29\text{k}$ $0.67\text{m}\pm 5.51\text{k}$ $0.79\text{m}\pm 17.16\text{k}$ $0.74\text{m}\pm 12.55\text{k}$ $0.70\text{m}\pm 16.92\text{k}$ $0.72\text{m}\pm 16.73\text{k}$ $0.72\text{m}\pm 16.60\text{k}$ $0.70\text{m}\pm 13.53\text{k}$
Table II: Quantitative Results on TUM-RGBD.
Cam Method Metric fr1/desk fr2/xyz fr3/office
Monocular MonoGS [9] ATE $4.93\pm 0.16$ $4.66\pm 0.13$ $3.35\pm 0.45$
FPS $1.87\pm 0.05$ $3.37\pm 0.06$ $2.26\pm 0.01$
IPF $84.07\pm 0.25$ $51.64\pm 0.26$ $60.5\pm 0.43$
PSNR $17.65\pm 0.40$ $15.56\pm 0.02$ $19.35\pm 0.31$
Points $\boldsymbol{26.64\text{k}\pm 1.58\text{k}}$ $43.59\text{k}\pm 2.09\text{k}$ $\boldsymbol{35.24\text{k}\pm 3.24\text{k}}$
Photo-SLAM [17] ATE $1.55\pm 0.06$ $0.63\pm 0.18$ $1.10\pm 0.70$
FPS $25.18\pm 0.30$ $25.83\pm 0.12$ $24.74\pm 0.25$
IPF $7.08\pm 0.08$ $6.66\pm 0.08$ $7.77\pm 0.20$
PSNR $19.69\pm 0.04$ $20.19\pm 0.52$ $18.32\pm 1.36$
Points $40.00\text{k}\pm 0.79\text{k}$ $0.10\text{m}\pm 7.50\text{k}$ $81.16\text{k}\pm 3.44\text{k}$
Ours (Photo-SLAM) ATE $1.55\pm 0.06$ $0.70\pm 0.08$ $0.57\pm 0.33$
FPS $24.95\pm 0.46$ $26.16\pm 0.12$ $25.03\pm 0.11$
IPF $17.88\pm 0.02$ $14.41\pm 0.26$ $16.06\pm 0.32$
PSNR $20.51\pm 0.08$ $21.54\pm 0.85$ $19.38\pm 1.47$
Points $38.65\text{k}\pm 1.82\text{k}$ $66.51\text{k}\pm 1.71\text{k}$ $51.71\text{k}\pm 3.46\text{k}$
RGBD Loopy-SLAM [28] ATE $3.93\pm 1.13$ $1.43\pm 0.16$ $4.65\pm 1.63$
FPS $0.23\pm 0.00$ $0.21\pm 0.00$ $0.20\pm 0.00$
IPF - - -
PSNR $13.66\pm 0.12$ $17.95\pm 0.41$ $17.43\pm 0.15$
Points - - -
SplaTAM [10] ATE $2.51\pm 0.01$ $0.50\pm 0.00$ $4.52\pm 0.21$
FPS $0.27\pm 0.01$ $0.03\pm 0.02$ $0.25\pm 0.00$
IPF $460.32\pm 0.00$ $460.88\pm 0.00$ $460.84\pm 0.00$
PSNR $21.03\pm 0.10$ $23.19\pm 0.13$ $20.10\pm 0.05$
Points $0.96\text{m}\pm 3.96\text{k}$ $6.36\text{m}\pm 81.37\text{k}$ $0.79\text{m}\pm 5.89\text{k}$
Gaussian-SLAM [11] ATE $2.74\pm 0.11$ $0.96\pm 0.44$ $8.42\pm 1.19$
FPS $0.57\pm 0.06$ $0.48\pm 0.03$ $0.59\pm 0.02$
IPF $309.37\pm 4.29$ $308.44\pm 0.04$ $310.66\pm 0.11$
PSNR $\boldsymbol{23.71\pm 0.10}$ $\boldsymbol{23.95\pm 0.39}$ $\boldsymbol{25.80\pm 0.09}$
Points $0.76\text{m}\pm 12.12\text{k}$ $0.69\text{m}\pm 26.07\text{k}$ $1.47\text{m}\pm 6.75\text{k}$
MonoGS [9] ATE $1.84\pm 0.09$ $1.71\pm 0.08$ $1.74\pm 0.10$
FPS $2.18\pm 0.02$ $3.23\pm 0.07$ $2.48\pm 0.03$
IPF $77.77\pm 0.06$ $51.23\pm 0.18$ $63.20\pm 0.06$
PSNR $19.00\pm 0.09$ $15.81\pm 0.03$ $19.11\pm 0.25$
Points $43.01\text{k}\pm 1.95\text{k}$ $\boldsymbol{37.20\text{k}\pm 4.78\text{k}}$ $52.67\text{k}\pm 2.00\text{k}$
Photo-SLAM [17] ATE $1.49\pm 0.03$ $0.32\pm 0.02$ $1.17\pm 0.34$
FPS $23.45\pm 0.18$ $23.44\pm 0.01$ $22.63\pm 0.22$
IPF $8.88\pm 0.14$ $7.68\pm 0.28$ $8.54\pm 0.26$
PSNR $19.98\pm 0.03$ $21.92\pm 0.42$ $22.18\pm 1.20$
Points $45.64\text{k}\pm 1.18\text{k}$ $68.68\text{k}\pm 10.00\text{k}$ $67.69\text{k}\pm 1.75\text{k}$
Ours (Photo-SLAM) ATE $1.52\pm 0.03$ $0.30\pm 0.01$ $0.90\pm 0.03$
FPS $23.06\pm 0.22$ $23.36\pm 0.07$ $22.78\pm 0.10$
IPF $20.60\pm 0.46$ $18.05\pm 0.31$ $17.66\pm 0.32$
PSNR $20.54\pm 0.06$ $22.75\pm 0.22$ $22.95\pm 0.79$
Points $38.65\text{k}\pm 0.76\text{k}$ $49.80\text{k}\pm 2.63\text{k}$ $71.33\text{k}\pm 6.79\text{k}$
GS-ICP-SLAM [16] ATE $3.26\pm 0.28$ $2.26\pm 0.04$ $3.07\pm 0.41$
FPS $30.00\pm 0.00$ $30.00\pm 0.00$ $30.00\pm 0.00$
IPF $6.10\pm 0.05$ $3.69\pm 0.05$ $3.96\pm 0.08$
PSNR $15.62\pm 0.07$ $18.43\pm 0.19$ $19.20\pm 0.05$
Points $0.53\text{m}\pm 6.82\text{k}$ $1.91\text{m}\pm 11.37\text{k}$ $2.09\text{m}\pm 21.04\text{k}$
Ours (GS-ICP-SLAM) ATE $3.92\pm 0.71$ $2.44\pm 0.06$ $4.11\pm 1.28$
FPS $30.00\pm 0.00$ $30.00\pm 0.00$ $30.00\pm 0.00$
IPF $20.02\pm 0.10$ $18.43\pm 0.15$ $12.17\pm 0.13$
PSNR $17.54\pm 0.07$ $21.35\pm 0.20$ $20.84\pm 0.06$
Points $0.18\text{m}\pm 3.65\text{k}$ $0.13\text{m}\pm 12.32\text{k}$ $0.34\text{m}\pm 19.24\text{k}$
Table III: Quantitative Results on VECtor.
Cam Method Metric corner-slow robot-normal corridors-dolly
Monocular Photo-SLAM [17] ATE $0.66\pm 0.01$ $2.20\pm 1.66$ $9.56\pm 6.08$
FPS $23.27\pm 0.21$ $21.90\pm 0.32$ $20.18\pm 0.26$
IPF $3.11\pm 0.03$ $3.37\pm 0.17$ $3.11\pm 0.03$
PSNR $24.63\pm 0.05$ $19.58\pm 0.18$ $15.31\pm 0.69$
Points $0.12\text{m}\pm 17.02\text{k}$ $0.16\text{m}\pm 72.38\text{k}$ $0.38\text{m}\pm 3.99\text{k}$
Ours (Photo-SLAM) ATE $0.68\pm 0.02$ $2.35\pm 1.17$ $10.06\pm 6.20$
FPS $21.56\pm 0.33$ $18.30\pm 1.20$ $18.00\pm 0.26$
IPF $7.69\pm 0.17$ $10.78\pm 0.68$ $11.11\pm 0.18$
PSNR $\boldsymbol{25.37\pm 0.12}$ $\boldsymbol{22.16\pm 1.46}$ $\boldsymbol{23.02\pm 5.67}$
Points $7.31\text{k}\pm 0.25\text{k}$ $8.24\text{k}\pm 2.06\text{k}$ $36.96\text{k}\pm 1.59\text{k}$
Stereo Photo-SLAM [17] ATE $1.15\pm 0.00$ $1.52\pm 0.00$ $11.91\pm 0.04$
FPS $20.43\pm 0.32$ $17.77\pm 0.31$ $19.31\pm 0.01$
IPF $1.68\pm 0.08$ $2.58\pm 0.04$ $2.76\pm 0.02$
PSNR $19.34\pm 0.02$ $16.59\pm 0.01$ $14.51\pm 0.34$
Points $38.98\text{k}\pm 4.29\text{k}$ $47.36\text{k}\pm 0.64\text{k}$ $0.24\text{m}\pm 2.92\text{k}$
Ours (Photo-SLAM) ATE $1.15\pm 0.00$ $1.52\pm 0.00$ $11.51\pm 0.23$
FPS $20.75\pm 0.37$ $14.64\pm 0.23$ $16.64\pm 0.83$
IPF $9.23\pm 0.02$ $12.24\pm 0.20$ $11.21\pm 0.16$
PSNR $19.56\pm 0.04$ $16.77\pm 0.05$ $19.34\pm 0.06$
Points $\boldsymbol{6.45\text{k}\pm 0.20\text{k}}$ $\boldsymbol{7.68\text{k}\pm 0.24\text{k}}$ $\boldsymbol{30.81\text{k}\pm 2.21\text{k}}$
@@
where $r_{k+1}^{0}$ is the initial optimization iteration count assigned to the new keyframe, and $l_{k+1}$ is its initial optimization loss value. We then select a keyframe $v^{\prime}$ randomly from the subset of keyframes with remaining iterations, defined as $\left\{v_{i}|r_{i}>0,\forall r_{i}\in\mathcal{R}_{k}\right\}$, to train the 3D Gaussians Map $\mathcal{G}$. Post-optimization, we decrement the optimization iteration count for the selected keyframe by one, adjusting $r^{\prime}$ to $r^{\prime}-1$, and also update the corresponding optimization loss value $l^{\prime}$.
@@
where $\overset{d_{k}}{\prod}(\cdot)$ donates top $d_{k}$ largest elements, $d_{k}=\max(1,\frac{k}{d})$, and $d$ is a hyperparameter. This method prioritizes keyframes with higher optimization loss values for the photorealistic rendering module, effectively tackling the long-tail optimization as demonstrated in LABEL:Fig:adaptive-opt.
III-C3 Opacity Regularization
In the typical application of 3DGS, the rendered loss $\mathcal{L}_{rendered}$ is utilized to refine the 3D Gaussian primitives [8]. To efficiently manage memory usage and model size, we have devised a strategy that encourages the elimination of Gaussians in areas where they do not contribute to the rendering process. Since the presence of a Gaussian is primarily indicated by its opacity $o$, we impose a regularization term $\mathcal{L}_{o}$ on this attribute. The complete formulation of our optimization loss $\mathcal{L}$ is as follows:
@@ IV Experiments
In this section, we present a comparative analysis of CaRtGS against state-of-the-art GS-SLAM systems [9, 10, 11, 17, 16] and Loopy-SLAM [28], a state-of-the-art NeRF-based SLAM system. This evaluation spans multiple scenarios, including those captured using monocular, RGB-D, and stereo cameras. Furthermore, we perform an ablation study to substantiate the efficacy of the novel techniques introduced in our approach.
@@ IV-A Setup
Dataset. We conducted evaluations on three distinct camera systems: monocular, RGB-D, and stereo. These assessments were carried out on three renowned datasets: Replica [29], TUM-RGBD [30], and VECtor [31]. Replica [29] is a high-quality reconstruction dataset at room and building scale, including high-resolution high-dynamic-range (HDR) textures. TUM-RGBD [30] is a well-known RGB-D dataset that contains color and depth images captured by a Microsoft Kinect sensor, along with the ground-truth trajectory obtained from a high-accuracy motion-capture system. VECtor [31] is a SLAM benchmark dataset that covers the full spectrum of motion dynamics, environmental complexities, and illumination conditions. To ensure data consistency, we employed a soft time synchronization to align the sensor data and ground truth with a precision of $\Delta t=0.08s$.
Implementation Detail. All experimental evaluations were conducted on a desktop with an Nvidia RTX 4090 GPU, an AMD Ryzen 9 7950X CPU, and 128 GB RAM. We retained most of the original hyperparameters from the 3DGS [8]. However, we densify every 500 iterations with a positional gradients threshold $\tau_{p}=0.001$ and remove the transparent Gaussians with a threshold $\epsilon_{\alpha}=0.02$. By default, we set $d=4$ and $\lambda_{o}=0.001$. On Replica, we use $r_{k+1}^{0}=8$, whileas $r_{k+1}^{0}=2$ on TUM-RGBD and VECtor.
Evaluation. We performed all experiments $5$ times to ensure statistical robustness and rendered original-resolution images for each estimated camera pose. To measure performance, we utilized the evo toolkit [^1] https://github.com/MichaelGrupp/evo and the torchmetrics toolkit [^2] https://github.com/Lightning-AI/torchmetrics. We recorded various performance indicators, including Absolute Trajectory Error (ATE) to assess the accuracy of localization, Peak Signal-to-Noise Ratio (PSNR) to assess the quality of the photorealistic renderings, and the number of 3D Gaussian points to assess the model size. To assess the sufficiency of the Gaussian primitives’ optimization, we introduced a metric known as Iterations Per Frame (IPF), defined as the ratio of total iterations to the total number of frames ($\text{IPF}=\frac{\text{Iterations}}{\text{Frames}}$). All performance indicators are reported in the format of mean $\pm$ standard deviation.